A basket contains 12 black marbles, 14 red marbles and 12 green marbles. What is the probability of picking a black marble first then a red marble as replacement?


1
Expert's answer
2021-07-12T17:35:36-0400

"Black\:marbles=12"

"Red\:marbles=14"

"\:Green\:marbles=12"

"Total\:marbles=38"


"P\left(picking\:black\:marble\right)=\frac{12}{38}"

"P\left(picking\:red\:marble\right)=\frac{14}{38}"

"\:P\left(picking\:green\:marble\right)=\frac{12}{38}"

Therefore, the probability of picking the first black marble and then the second red marble with replacement is:

"P\left(black\:marble\:then\:red\:marble\right)=\frac{12}{38}\times \frac{14}{38}=\frac{42}{361}"